# UNB StepUP: A footStep database for gait analysis and recognition using Underfoot Pressure

### v2

@@ Abstract
Gait refers to the patterns of limb movement generated during walking, which are unique to each individual due to both physical and behavioural traits. Walking patterns have been widely studied in biometrics, biomechanics, sports, and rehabilitation. While traditional methods rely on video and motion capture, advances in underfoot pressure sensing technology now offer deeper insights into gait. However, underfoot pressures during walking remain underexplored due to the lack of large, publicly accessible datasets. To address this, the UNB StepUP database was created, featuring gait pressure data collected with high-resolution pressure sensing tiles (4 sensors/cm2, 1.2m by 3.6m). Its first release, UNB StepUP-P150, includes over 200,000 footsteps from 150 individuals across various walking speeds (preferred, slow-to-stop, fast, and slow) and footwear types (barefoot, standard shoes, and two personal shoes). As the largest and most comprehensive dataset of its kind, it supports biometric gait recognition while presenting new research opportunities in biomechanics and deep learning. The UNB StepUP-P150 dataset sets a new benchmark for pressure-based gait analysis and recognition.
@@
Gait analysis has long been a critical area of study within biometrics, biomechanics, sports, and rehabilitation. The unique patterns of human walking provide valuable insight into an individual’s identity, health status, athletic performance, and even potential underlying medical conditions. Although traditional methods of gait analysis have relied heavily on video-based systems and motion capture technologies, recent advances in sensor technology have opened new avenues for capturing and analyzing gait patterns. Among these, underfoot pressure measurement has appeared as a promising avenue due to its potential to provide detailed insights into footstep patterns, including the pressure exerted at different points under the foot and temporal changes in these patterns throughout the gait cycle. Despite its potential, this modality remains under-researched [1], especially with regard to the development of large, publicly available datasets, thereby limiting the exploration of modern advanced deep learning algorithms for gait analysis and recognition.
Despite the availability of many other gait databases and datasets, the majority concentrate on video-based systems, motion capture technologies, and wearable devices [2, 3, 4, 5]. Databases that include floor sensor technology generally collect data from force plates, as seen in the GaitRec [6], Gutenberg [7], and ForceID A [8] datasets, which provide more constrained and less comprehensive information than the complex spatio-temporal data offered by emerging high-resolution pressure sensors. Conversely, databases that emphasize high-resolution underfoot pressure data, such as the CASIA-D [9], SFootBD [10], CAD WALK [11], and UoM-Gait-69 [12] datasets, usually address a limited scope, often with little or no consideration of covariates that may confound performance, and generally involve small sample sizes for subjects and walking trials, thereby under-representing real-world scenarios. To fill this gap, UNB StepUP: A footStep database for gait analysis and recognition using Underfoot Pressure, was developed. The purpose of creating the UNB StepUp database is to compile a diverse range of datasets that include detailed foot pressure measurements, explore various covariate factors, involve a large number of participants, and encompass numerous walking trials. The initial StepUP experiments include three separate studies, each characterized by unique covariates: (1) walking speed and footwear, (2) different sessions (returning from study 1) and load carriage (e.g. carrying objects of different shapes and weights), and (3) walking environment, clothing, viewing angle (in terms of sensor orientation and placement relative to the subject), and dual-task gait (which involves walking while simultaneously engaging in a secondary task). In addition to pressure data, concurrent multi-angle video recordings and digital footprints and shoeprints were also recorded and will be shared as separate complementary and matched datasets in the future.
@@
This initial paper on the UNB StepUP database presents the UNB StepUP-P150 dataset, derived from the first of the three studies. The designation P150 refers to pressure data collected from 150 participants. The dataset includes more than 200,000 footsteps from 150 individuals, covering a range of walking speeds (self-paced, self-paced with a stop at the end, fast, and slow) and different types of footwear (barefoot, standard shoes, and two types of personal shoes), making it the most comprehensive openly available collection of its kind (Table 1). The primary method for data collection involved high-resolution pressure tiles (a walkway of dimensions 1.2 meters by 3.6 meters, containing 240 by 720 pressure sensors) to capture the pressure distribution of each footstep during natural gait. A total of 24 minutes of walking data were collected from each participant, spread across 16 different walking conditions (4 walking speeds $\times$ 4 types of footwear) at 90 seconds per condition, resulting in approximately 1,400 steps per individual. Compared to the largest previously published footstep database, SFootBD, which recorded nearly 20,000 footsteps from 127 individuals [13], with only 5 subjects contributing more than 1,000 footsteps each, this initial UNB StepUP-P150 dataset surpasses SFootBD by more than tenfold, making it the largest footstep dataset to date.
@@
The UNB StepUP-P150 dataset consists of raw trial-by-trial pressure data along with preprocessed data segmented by each footstep. This dual format allows users to either work directly with the raw data or leverage the pre-processed data for rapid prototyping and analysis. To ensure its quality and consistency, the foot pressure data was subjected to thorough quality control and pre-processing, such as footstep segmentation, foot alignment, and temporal normalization. This information is complemented by contextual metadata detailing the demographics of the participants and the conditions under which the data were collected (including variables such as walking speed, footwear types, foot side, and incomplete footsteps), providing a comprehensive and fully annotated gait resource for researchers. In addition, video data were used for visual examination of gait events and patterns, helping to verify uncertain labels generated by automatic algorithms and providing additional context during the preprocessing and manual inspection of pressure data. The finalized dataset is designed for ease of use by researchers and is compatible with various analytical tools, including support of both NPZ (Python) and MAT (MATLAB) file formats.
@@
The primary aim of the UNB StepUP database is to support the advancement of gait analysis and gait recognition, motivated by applications in biometric systems. To mitigate demographic biases within the dataset, the study includes participants of various ages, sexes/genders, races/ethnicities, and body sizes. This ensures that the dataset accurately captures the broad variability inherent in human gait, making it both diverse and representative of the population. This diversity is critical for developing generalizable models that can perform accurately across different population groups. In developing this gait dataset with attention to demographic bias in biometrics, the UNB StepUP-P150 dataset opens up numerous research opportunities in gait analysis and recognition, extending beyond biometric uses. Researchers in fields ranging from biomechanics to machine learning and deep learning can take advantage of the UNB StepUP-P150 to train and test new models, explore the relationship between underfoot pressure and other gait metrics, or investigate how various factors, such as sex, age, walking speed, or footwear, affect gait patterns. Furthermore, the dataset can be used as a benchmark for comparing different gait analysis and recognition techniques, thereby contributing to the standardization of methodologies in the field. The UNB StepUP-P150 dataset offers a level of detail and comprehensiveness that is unmatched by existing underfoot pressure-based gait datasets, making it possible to explore new avenues of research in gait analysis, footstep recognition, and related areas.
@@ Methods
A total of 180 individuals participated in the study, which took place over a period of 18 months at the University of New Brunswick (UNB) in Fredericton, New Brunswick, Canada. The participants consisted of students, staff, and faculty from the university, and community members who were recruited via word-of-mouth and promotional materials circulated in the local community. The general inclusion criteria were: (1) at least 18 years of age; (2) able to walk comfortably without assistive devices for 90 seconds at a time with different walking speeds and footwear conditions (24 minutes total); and (3) able and comfortable to balance unassisted, on both feet and on one foot, for 30 seconds at a time with different footwear conditions (6 minutes total). Before collecting data, all participants provided their written informed consent to participate per the Declaration of Helsinki. The study protocol was reviewed and approved by the University of New Brunswick’s Research Ethics Board (REB 2022-132).
@@
The following factors led to participant exclusion from the final dataset to maintain its quality and integrity: (1) experimental deviations, e.g. unintended behaviour during one or more trials ($N=2$); (2) missing data, e.g. incomplete protocols or corrupted files ($N=4$); (3) hardware malfunction, e.g. disconnected sensors significantly affecting recordings ($N=15$), and (4) no accompanying video data to confirm annotations, as participants did not consent to video collection ($N=9$). Consequently, 30 participants were removed from the final dataset, resulting in a total of 150 subjects being included. Although not made publicly available as part of the UNB StepUP-P150 dataset, portions of the data from these 30 excluded experiments may be used to support future footstep biometric competitions as independent test samples.
@@
Table 2 provides a summary of the demographic and anthropometric data for the 150 participants in the study. The UNB StepUP-P150 dataset features individuals with an average age of 34 years, spanning from 19 to 91 years. This cohort comprises 74 males and 76 females. The dataset maintains an approximately balanced sex/gender distribution and includes both younger and older adults, with the age distribution for participants by sex illustrated in Fig. 1(a). Most of the participants identify as White ($N=106$, $71\%$), with additional identifications as Asian ($N=36$, $24\%$; 15 Middle Eastern, 10 South Asian, 11 East/Southeast Asian), Other/Multiple ($N=6$, $4\%$; 1 Black, 5 multi-ethnic), and Unknown/Not Specified ($N=2$, $1\%$). This reflects an over-representation of visible minority groups relative to the official population distribution in the local community; the proportion of total visible minorities in the database is 28%, as opposed to the 14% reported by Statistics Canada [14]. The dataset also includes participants with a variety of body types and dimensions, including height ranging from 151 to 196 cm, weight between 46 and 148 kg, BMI spanning 17 to 39 kg/m2, foot length measuring 20 to 30 cm, foot width between 7 and 11 cm, and UK shoe sizes from 4 to 12.5 (Fig. 1(b)). Figure 2 shows the distribution of participants’ body sizes and foot measurements, categorized by both sex and ethnicity/race.
@@
Pressure data from footsteps were collected using a specialized runway (Fig. 3), which featured a $2\times 6$ grid of commercial pressure sensing tiles developed by Stepscan Technologies Inc. The runway has a total active recording area of 1.2 m $\times$ 3.6 m, allowing for the collection of multiple consecutive steps (typically 4-6 steps). Each modular 60 cm $\times$ 60 cm tile is made up of a $120\times 120$ sensor grid, corresponding to a spatial resolution of 4 sensors/cm2. The set of 172,800 ($240\times 720$) embedded sensors can detect pressures within a range of 1,510 kPa, have a threshold sensitivity of 10 kPa, feature a resolution below 10 kPa, and were sampled at a rate of 100 Hz during the recording process. To increase system durability, particularly in areas experiencing high foot traffic and the use of outdoor footwear, as anticipated in the practical application of this study, the manufacturer tailored the tiles with a 6 mm protective layer made of Ramflex double-layered vulcanized rubber flooring, instead of the usual 2 mm thickness commonly applied in clinical usage of this technology in healthcare applications. These tiles were installed in a laboratory reserved for data collection and stayed there throughout the entire study. A non-instrumented wooden platform was also constructed around the tile grid, extending 1.4 meters at each end of the runway. The platform was flush with the tiles to allow natural walking behaviour, eliminating the need to step up or down and providing room for the participants to turn during the experiment (Fig. 3). It should be noted that the pressure-sensitive tiles are pre-calibrated by the manufacturer (Stepscan Technology Inc.), and did not require on-site calibration. Each tile comes with factory-set parameters that align the signal characteristics of each physical pressure sensor with a common logical signal domain. The companion data collection application automatically implements these corrections, ensuring that they are reflected in all data within this dataset. Therefore, no further calibration efforts were undertaken by the research team.
@@
In addition to pressure-sensing tiles, seven RGB video cameras (QCN8068BA, Q-See, USA) with a resolution of 1440p ($2560\times 1440$) and a frame rate of 20 fps were strategically installed around the room. These cameras were positioned to record participants’ gait from various angles: approximately 0°, 45°, 90°, 135°, 225°, 270°, and 315° with respect to the major axis of the tile grid (Fig. 3). Participants in the UNB StepUP-P150 dataset consented to have their sessions videotaped, producing seven separate video recordings for each pressure data capture. In addition, a flatbed scanner (OpticSlim 1180, Plustek, USA) with an optical resolution of 1200 dpi, 48-bit color input, and a scanning area of 29.7 cm $\times$ 43.18 cm was used to acquire digital scans of the participant’s two pairs of personal footwear that they brought to their session. The pressure and video recordings were captured simultaneously on two networked desktop computers, connected via the Secure Shell (SSH) protocol, and time-synchronized using the Network Time Protocol (NTP). Each recording was time-stamped to enable precise offline synchronization of the data from the two sensor types.
@@
Figure 4 provides a summary of the 90-minute experimental protocol. The initial preparation session takes around 30 minutes, followed by balance and walking sessions that last for about an hour. Before starting the walking trials, three separate 30-second balance tests were performed for each type of footwear. The walking experiment then involved sixteen 90-second walking trials, featuring four different footwear conditions, each executed at four distinct walking speeds. As a result, there were seven recorded trials for each type of footwear (consisting of three 30-second balance tasks and four 90-second walking tasks), resulting in 28 trials overall and a total recording duration of 30 minutes. Participants were allowed to take breaks and rest between trials as needed, and whenever there was a need to change shoes, a minimum of 2 minutes was allocated for breaks. To minimize the potential impact of any fatigue that may affect recorded gait patterns, the order of footwear and walking speed conditions was randomized for each participant (see Fig. 4 for an example sequence). The subsequent sections provide a more detailed explanation of different tasks and conditions.
@@
Before the scheduled collection session, the participants were informed that the session would consist of a series of walking and balance tasks, and they were asked to bring along two pairs of their own shoes for the experiment. Upon their arrival, participants received an overview of the experiment’s procedure and provided their signatures on the informed consent form (UNB REB 2022-132). The participants subsequently filled out a form to provide essential demographic details, including sex, gender, birth year, race or ethnicity (options included Aboriginal, Black, East/Southeast Asian, Latino, Middle Eastern, South Asian, White, or Other), along with pertinent physical information that could influence their gait patterns, such as any recent injuries and which leg is dominant, determined by answering “Which foot would you normally use to kick a stationary ball straight in front of you?”[15]. For the precise determination of the anthropometric parameters, the experimenter helped the participants measure their height, body weight, and foot size. Body weight was measured using a Withings body smart scale, height with a Seca 213 portable stadiometer, and foot length and width with an adult Brannock device. In addition, digital shoe sole scans were obtained from the two pairs of participant-provided footwear, with the sizes, brands, and descriptions of each shoe documented. The participants were then equipped with a pair of standard shoes supplied by the research team, specifically Adidas Grand Court 2.0 unisex sneakers, available in sizes from US Men’s size 4 (Women’s size 5) to US Men’s size 13 (Women’s size 14), inclusive of half sizes. This is equivalent to UK sizes 3.5 to 12.5. Participants were instructed to select the size that offered the most comfort and to take a few steps in the shoes to verify fit. Their preferred sizes were thereafter documented.
@@
For the walking trials, the participants were instructed to walk back and forth across the tiles naturally for 90 seconds, specifically along the longer 3.6 m grid direction, making use of the non-instrumented portion of the platform to turn around (Fig. 3). The four walking speed conditions were as follows:
@@
The average self-selected preferred (moderate) pace was 1.12 m/s, the average fast speed was 1.45 m/s (which is $26\%$ faster) and the average slow speed was 0.83 m/s (which is $29\%$ slower) (Fig. 5). It should be noted that during the trials, the participants had the freedom to choose how they navigated their walking paths, including making 180-degree turns between passes in any way they preferred. However, the experimenters advised the participants to change their turning directions and/or pause on the non-instrumented landing when necessary to reduce the possibility of dizziness.
@@
The trials were completed with four different footwear conditions as described below:
@@
  • Standard Shoe (ST): a common pair of flat-soled casual sneakers (Adidas Grand Court 2.0).
  • Personal Shoes (P1 and P2): two pairs of personal shoes frequently worn in everyday situations. The personal shoes included a wide range of different shoe types, which have been broadly categorized as athletic sneakers (e.g., road running shoes, indoor trainers), casual sneakers (e.g., skate shoes, court-inspired sneakers, high-top leather sneakers), sandals (e.g., slides, flip-flops, ankle-strap sandals), flat canvas shoes (e.g., lace-up or slip-on shoes with canvas upper), boots (e.g., winter boots, steel-toe work boots, Chelsea boots), business/dress shoes (e.g., high heels, ballet flats, oxfords), hiking/trail shoes (e.g., trail running shoes, outdoor walking shoes), and other (e.g., foam clogs, minimal shoes). Figure 6 presents the breakdown of shoes brought in by the participants, grouped into these eight categories. The most commonly brought shoes were athletic sneakers ($N=114$), followed by casual sneakers ($N=53$) and sandals/flip flops ($N=35$). A smaller number of shoes with unique pressure patterns, such as high heels ($N=3$) and steel or composite toe work boots ($N=5$), were also acquired.
Figure 7 shows the time series of pressure images for an individual taking a single footstep, highlighting changes in the percentage of stance phase across four different footwear conditions, whereas Figure 8 illustrates variations in peak pressure profiles across different participants and shoe types.
Additional Covariates
Although not explicitly elicited or controlled in the first data collection protocol, several additional covariates may be present in the UNB StepUP-P150 dataset and introduce complexity from the perspective of gait biometrics and gait recognition. As noted previously, these covariate factors are controlled and examined in the second and third protocols.
  • Viewing angle: Changes in a sensor’s “view,” including its alignment and positioning relative to the subject, pose a significant challenge, especially in vision and accelerometry modalities. In pressure-based gait, the alignment of the foot relative to the sensing platform may be considered as the viewing angle. This concern has not been extensively addressed in earlier force-/pressure-based gait studies, primarily due to the constraints of force plate and pressure mat implementations, which allow movement along a single, narrow straight line. Consequently, the foot angles across different trials and subjects are fairly consistent, showing minimal variation with small angles relative to the platform’s main axis. In addition, earlier studies frequently advise rotating or aligning foot images to standardize the pressure feature representations. In the present study, although participants usually walked in a straight line, the expanded platform area allowed them to choose their routes. Consequently, there were occasions where the path was slightly curved or diagonal.
  • Clothing: Changes in an individual’s attire can pose a significant challenge, particularly within vision and acoustic modalities, yet they likely have minimal impact on pressure-based gait recognition. This study recommended that participants wear comfortable attire for walking; however, there was no strict regulation or control over their clothing choices. It should be mentioned that none of the 150 participants changed their attire (i.e., other than the footwear changes described above) during the 90-minute experiment, resulting in no variation within individual subjects.
  • Time elapsed: While the impact of elapsed time on gait recognition may seem significant, it is primarily attributed to related covariates. The initial and final footsteps collected for each participant in this dataset occur within one hour, with variations in either footwear, walking speed, or both, as determined by a randomized sequence.
  • Load carriage, dual-task gait, and walking environment: In the initial experiment, participants are not allowed to carry additional weight such as a backpack, engage in secondary tasks like using a mobile phone, or encounter any changes in the walking environment or setting (such as obstacles in the path).

### v3

# A dataset of high-resolution plantar pressures for gait analysis across varying footwear and walking speeds

@@ Abstract
Gait refers to the patterns of limb movement generated during walking, which are unique to each individual due to both physical and behavioural traits. Walking patterns have been widely studied in biometrics, biomechanics, sports, and rehabilitation. While traditional methods rely on video and motion capture, advances in plantar pressure sensing technology now offer deeper insights into gait. However, underfoot pressures during walking remain underexplored due to the lack of large, publicly accessible datasets. To address this, we introduce the UNB StepUP-P150 dataset: a footStep database for gait analysis and recognition using Underfoot Pressure, including data from 150 individuals. This dataset comprises high-resolution plantar pressure data (4 sensors/cm2) collected using a 1.2m by 3.6m pressure-sensing walkway. It contains over 200,000 footsteps from participants walking with various speeds (preferred, slow-to-stop, fast, and slow) and footwear conditions (barefoot, standard shoes, and two personal shoes), supporting advancements in biometric gait recognition and presenting new research opportunities in biomechanics and deep learning. UNB StepUP-P150 establishes a new benchmark for plantar pressure-based gait analysis and recognition.
@@
Gait analysis has long been a critical area of study within biometrics, biomechanics, sports, and rehabilitation. The unique patterns of human walking provide valuable insight into an individual’s identity, health status, athletic performance, and even potential underlying medical conditions. Although traditional methods of gait analysis have relied heavily on video-based systems and motion capture technologies, recent advances in sensor technology have opened new avenues for capturing and analyzing gait patterns. Among these, plantar pressure measurement has appeared as a promising avenue due to its potential to provide detailed insights into footstep patterns, including the pressure exerted at different points under the foot and temporal changes in these patterns throughout the gait cycle. Despite its potential, this modality remains under-researched [1], especially with regard to the development of large, publicly available datasets, thereby limiting the exploration of modern advanced deep learning algorithms for gait analysis and recognition.
@@
Despite the availability of many other gait databases and datasets, the majority use video-based systems, motion capture technologies, and wearable devices [2, 3, 4, 5]. Databases that use floor sensor technology generally collect data from force plates, as seen in the GaitRec [6], Gutenberg [7], and ForceID A [8] datasets, which provide more constrained and less comprehensive information than the complex spatio-temporal data offered by emerging high-resolution pressure sensors. Conversely, databases that emphasize high-resolution underfoot pressure data, such as the CASIA-D [9], SFootBD [10], CAD WALK [11], and UoM-Gait-69 [12] datasets, usually address a limited scope, often with little or no consideration of covariates that may confound performance, and generally involve small sample sizes for subjects and walking trials, thereby under-representing real-world scenarios (see Table 1). To fill this gap, UNB StepUP-P150: A footStep dataset for gait analysis and recognition using Underfoot Pressure, was developed. The designation P150 refers to pressure data collected from 150 participants. The dataset includes more than 200,000 footsteps from these 150 individuals, covering a range of walking speeds (self-paced, self-paced with a stop at the end, fast, and slow) and different types of footwear (barefoot, standard shoes, and two types of personal shoes).The primary method for data collection involved high-resolution pressure tiles (a walkway of dimensions 1.2 meters by 3.6 meters, containing 240 by 720 pressure sensors) to capture the pressure distribution of each footstep during natural gait. A total of 24 minutes of walking data were collected from each participant, spread across 16 different walking conditions (4 walking speeds $\times$ 4 types of footwear) at 90 seconds per condition, resulting in approximately 1,400 steps per individual. Compared to the largest previously published footstep database, SFootBD, which recorded nearly 20,000 footsteps from 127 individuals [13], with only 5 subjects contributing more than 1,000 footsteps each, the UNB StepUP-P150 dataset surpasses SFootBD by more than tenfold, making it the largest footstep dataset to date.
@@
The UNB StepUP-P150 dataset provides both raw trial-by-trial pressure data and preprocessed data segmented by each footstep. This dual format allows users to either work directly with the raw data or leverage the preprocessed data for rapid prototyping and analysis. To ensure its quality and consistency, the foot pressure data was subjected to thorough quality control and preprocessing, such as footstep segmentation, foot alignment, and temporal normalization. Data records are complemented by contextual metadata detailing the demographics of the participants and the conditions under which the each footstep was collected (including variables such as walking speed, footwear types, foot side, and incomplete footsteps), providing a comprehensive and fully annotated gait resource for researchers. In addition, video data were used for visual examination of gait events and patterns, helping to verify uncertain labels generated by automatic algorithms and providing additional context during the preprocessing and manual inspection of pressure data. The finalized dataset is designed for ease of use by researchers and is compatible with various analytical tools, including support of both NPZ (Python) and MAT (MATLAB) file formats.
@@
The primary aim of this work is to support the advancement of gait analysis and gait recognition, motivated by applications in biometric systems. To mitigate demographic biases within the dataset, the study includes participants of various ages, sexes/genders, races/ethnicities, and body sizes. This ensures that the dataset accurately captures the broad variability inherent in human gait, making it both diverse and representative of the population. In developing this gait dataset with attention to demographic bias in biometrics, the UNB StepUP-P150 dataset opens up numerous research opportunities in gait analysis and recognition, extending beyond biometric uses. Researchers in fields ranging from biomechanics to machine learning and deep learning can take advantage of the UNB StepUP-P150 to train and test new models, explore the relationship between underfoot pressure and other gait metrics, or investigate how various factors, such as sex, age, walking speed, or footwear, affect gait patterns. Furthermore, the dataset can be used as a benchmark for comparing different gait analysis and recognition techniques, thereby contributing to the standardization of methodologies in the field. The UNB StepUP-P150 dataset offers a level of detail and comprehensiveness that is unmatched by existing underfoot pressure-based gait datasets, making it possible to explore new avenues of research in gait analysis, footstep recognition, and related areas. In addition to the primary biometric focus, UNB StepUP-P150 can support research topics ranging from:
  • statistical models of normative walking gait based on foot pressure patterns,
  • differences in pressure-based gait patterns between demographic subgroups, categorized by factors like gender and age,
  • differences in pressure patterns resulting from external factors such as different walking speeds and types of footwear,
  • novel machine learning and deep learning models for gait recognition across various classification challenges,
  • novel approaches for the segmentation, alignment, and/or registration of plantar pressure images, and
  • development and evaluation of state-of-the-art techniques for gait analysis and gait recognition.
In summary, UNB StepUP-P150 establishes a new benchmark for plantar pressure-based gait research.
@@ Methods
A total of 180 individuals participated in the study, which took place over a period of 18 months at the University of New Brunswick (UNB) in Fredericton, New Brunswick, Canada. The participants consisted of students, staff, and faculty from the university, and community members who were recruited via word-of-mouth and promotional materials circulated to local businesses and clubs. The general inclusion criteria were: (1) at least 18 years of age; (2) able to walk comfortably without assistive devices for 90 seconds at a time with different walking speeds and footwear conditions (24 minutes total); and (3) able and comfortable to balance unassisted, on both feet and on one foot, for 30 seconds at a time with different footwear conditions (6 minutes total). Before collecting data, all participants provided their written informed consent for their participation and for the inclusion of their data in the public dataset release. The study protocol was reviewed and approved by the University of New Brunswick’s Research Ethics Board (REB 2022-132), in accordance with the Declaration of Helsinki.
@@
The following factors led to participant exclusion from the final dataset to maintain its quality and integrity: (1) experimental deviations, e.g. unintended behaviour during one or more trials ($N=2$); (2) missing data, e.g. incomplete protocols or corrupted files ($N=4$); (3) hardware malfunction, e.g. disconnected sensors significantly affecting recordings ($N=15$), and (4) no accompanying video data to confirm annotations, as participants did not consent to video collection ($N=9$). Consequently, 30 participants were removed from the final dataset, resulting in a total of 150 subjects being included. None of the participants reported having a neurological condition (e.g., stroke or Parkinson’s), a concussion within six months of recording, or pregnancy. Moreover, although not made publicly available as part of the UNB StepUP-P150 dataset, portions of the data from the 30 excluded experiments will be used to support future footstep biometric competitions as independent test samples.
@@
Table 2 provides a summary of the demographic and anthropometric data for the 150 participants in the study. The dataset includes 74 males and 76 females with an average age of 34 years, spanning 19 to 91 years. The dataset maintains an approximately balanced sex/gender distribution and includes both younger and older adults, with the age distribution for participants by sex illustrated in Fig. 1(a). Most of the participants identify as White ($N=106$, $71\%$), with additional identifications as Asian ($N=36$, $24\%$; 15 Middle Eastern, 10 South Asian, 11 East/Southeast Asian), Other/Multiple ($N=6$, $4\%$; 1 Black, 5 multi-ethnic), and Unknown/Not Specified ($N=2$, $1\%$). This reflects an over-representation of visible minority groups relative to the official population distribution in the local community; the proportion of total visible minorities in the database is 28%, as opposed to the 14% reported by Statistics Canada [14]. The dataset also includes participants with a variety of body types and dimensions, including height ranging from 151 to 196 cm, weight between 46 and 148 kg, BMI spanning 17 to 39 kg/m2, foot length measuring 20 to 30 cm, foot width between 7 and 11 cm, and UK shoe sizes from 4 to 12.5 (Fig. 1(b)). Figure 2 shows the distribution of participants’ body sizes and foot measurements, categorized by both sex and ethnicity/race.
@@
Pressure data from footsteps were collected using a specialized runway (Fig. 3), which featured a $2\times 6$ grid of commercial pressure sensing tiles developed by Stepscan Technologies Inc. The runway has a total active recording area of 1.2 m $\times$ 3.6 m, allowing for the collection of multiple consecutive steps (typically 4-6 steps). Each modular 60 cm $\times$ 60 cm tile is made up of a $120\times 120$ sensor grid, corresponding to a spatial resolution of 4 sensors/cm2. Collectively, the $2\times 6$ tile grid comprises 172,800 piezoresistive sensors (i.e., $240\times 720$), each with a threshold sensitivity of 10 kPa, sensing range of 1,510kPa (10 kPa resolution), and 100 Hz sampling rate. To increase system durability, particularly in areas experiencing high foot traffic and the use of outdoor footwear, as anticipated in the practical application of this study, the manufacturer tailored the tiles with a 6 mm protective layer made of Ramflex double-layered vulcanized rubber flooring, instead of the usual 2 mm thickness commonly applied in clinical usage of this technology for healthcare applications. These tiles were installed in a laboratory reserved for data collection and stayed there throughout the entire study. A non-instrumented wooden platform was also constructed around the tile grid, extending 1.4 meters at each end of the runway. The platform was flush with the tiles to allow natural walking behaviour, eliminating the need to step up or down and providing room for the participants to turn during the experiment. It should be noted that the pressure-sensitive tiles are pre-calibrated by the manufacturer (Stepscan Technologies Inc.), and did not require on-site calibration. Each tile comes with factory-set parameters that align the signal characteristics of each physical pressure sensor with a common logical signal domain. The companion data collection application automatically implements these corrections, ensuring that they are reflected in all data within this dataset. Therefore, no further calibration efforts were undertaken by the research team.
@@
In addition to pressure-sensing tiles, seven RGB video cameras (QCN8068BA, Q-See, USA) with a resolution of 1080p ($1920\times 1080$) and a frame rate of 20 fps were strategically installed around the room. These cameras were positioned to record participants’ gait from various angles: approximately 0°, 45°, 90°, 135°, 225°, 270°, and 315° with respect to the major axis of the tile grid (Fig. 3(a)). Participants in the UNB StepUP-P150 dataset consented to have their sessions videotaped, producing seven separate video recordings for each pressure data capture. In addition, a flatbed scanner (OpticSlim 1180, Plustek, USA) with an optical resolution of 300 dpi, 48-bit color input, and a scanning area of 29.7 cm $\times$ 43.18 cm was used to acquire digital scans of the participant’s two pairs of personal footwear that they brought to their session. The pressure and video recordings were captured simultaneously on two networked desktop computers, connected via the Secure Shell (SSH) protocol, and time-synchronized using the Network Time Protocol (NTP). Each recording was time-stamped to enable precise offline synchronization of the data from the two sensor types. Although video recordings and digital shoe scans are not included in the StepUP-P150 release, these were used internally by the research team to manually confirm metadata and pressure data integrity (see Technical Validation, below).
@@
Figure 4 provides a summary of the 90-minute experimental protocol. The initial preparation session took around 30 minutes, followed by about an hour of balance and walking sessions. Before starting the walking trials, three separate 30-second balance tests were performed for each type of footwear. The walking experiment then involved sixteen 90-second walking trials, featuring four different footwear conditions, each executed at four distinct walking speeds. As a result, there were seven recorded trials for each type of footwear, resulting in 28 trials overall and a total recording duration of 30 minutes. Participants were allowed to take breaks and rest between trials as needed, and whenever there was a need to change shoes, a minimum of 2 minutes was allocated for rest. To minimize the potential impact of fatigue or ordering effects on the recorded gait patterns, the order of footwear and walking speed conditions was randomized for each participant (see Fig. 4 for an example sequence). The subsequent sections provide a more detailed explanation of different tasks and conditions.
@@
Before the scheduled collection session, the participants were informed that the session would consist of a series of walking and balance tasks, and they were asked to bring along two pairs of their own shoes for the experiment. Upon their arrival, participants received an overview of the experimental procedure. They then filled out a form to provide essential demographic details, including sex, gender, birth year, race or ethnicity (options included Aboriginal, Black, East/Southeast Asian, Latino, Middle Eastern, South Asian, White, or Other), along with pertinent physical information that could influence their gait patterns, such as any recent injuries and leg dominance, determined by answering “Which foot would you normally use to kick a stationary ball straight in front of you?”[15]. To obtain precise anthropometric measurements, the experimenter assisted each participant in recording their height, body weight, and foot size. Body weight was measured using a Withings body smart scale, height with a Seca 213 portable stadiometer, and foot length and width with an adult Brannock device. In addition, digital shoe sole scans were obtained from the two pairs of participant-provided footwear, with the sizes, brands, and descriptions of each shoe documented. The participants were then fitted with a pair of standard shoes supplied by the research team, specifically Adidas Grand Court 2.0 unisex sneakers, available in sizes from US Men’s size 4 (Women’s size 5) to US Men’s size 13 (Women’s size 14), inclusive of half sizes. This is equivalent to UK sizes 3.5 to 12.5. Participants were instructed to select the size that offered the most comfort and to take a few steps in the shoes to verify fit. Their preferred sizes were thereafter documented.
@@
For the walking trials, the participants were instructed to walk back and forth across the tiles naturally for 90 seconds along the longer 3.6 m grid direction, making use of the non-instrumented portion of the platform to turn around (Fig. 3). The four walking speed conditions were:
@@
The average self-selected preferred (moderate) pace was 1.12 m/s, the average fast speed was 1.45 m/s (which is $26\%$ faster) and the average slow speed was 0.83 m/s (which is $29\%$ slower) (see Fig. 5). It should be noted that during the trials, the participants had the freedom to choose any walking path and to make the 180-degree turns between passes in any way they preferred. However, the experimenters advised the participants to change their turning directions and/or pause on the non-instrumented landing when necessary to reduce the possibility of dizziness.
@@
The trials were completed with four different footwear conditions:
@@
  • Standard Shoe (ST): a common pair of flat-soled casual sneakers (Adidas Grand Court 2.0) provided by the research team.
  • Personal Shoes (P1 and P2): two pairs of personal shoes frequently worn by the participants. The personal shoes included a wide range of different shoe types, which have been broadly categorized as athletic sneakers (e.g., road running shoes, indoor trainers), casual sneakers (e.g., skate shoes, court-inspired sneakers, high-top leather sneakers), sandals (e.g., slides, flip-flops, ankle-strap sandals), flat canvas shoes (e.g., lace-up or slip-on shoes with canvas upper), boots (e.g., winter boots, steel-toe work boots, Chelsea boots), business/dress shoes (e.g., high heels, ballet flats, oxfords), hiking/trail shoes (e.g., trail running shoes, outdoor walking shoes), and other (e.g., foam clogs, minimal shoes). Figure 6 presents the breakdown of shoes brought in by the participants, grouped into these eight categories. The most commonly brought shoes were athletic sneakers ($N=114$), followed by casual sneakers ($N=53$) and sandals/flip flops ($N=35$). A smaller number of shoes with unique pressure patterns, such as high heels ($N=3$) and steel or composite toe work boots ($N=5$), were also included.
Figure 7 shows the time series of pressure images for an individual taking a single footstep, highlighting changes across the stance phase in four different footwear conditions, whereas Figure 8 illustrates variations in peak pressure profiles across different participants and shoe types.